# A Multi-Fidelity Control Variate Approach for Policy Gradient Estimation

### v3

@@ Abstract
Many reinforcement learning (RL) algorithms are impractical for deployment in operational systems or for training with computationally expensive high-fidelity simulations, as they require large amounts of data. Meanwhile, low-fidelity simulators—such as reduced-order models, heuristic reward functions, or generative world models—can cheaply provide useful data for RL training, even if they are too coarse for direct sim-to-real transfer. We propose multi-fidelity policy gradients (MFPGs), an RL framework that mixes a small amount of data from the target environment with a control variate formed from a large volume of low-fidelity simulation data to construct an unbiased, variance-reduced estimator for on-policy policy gradients. We instantiate the framework by developing a practical, multi-fidelity variant of the classical REINFORCE algorithm. We show that under standard assumptions, the MFPG estimator guarantees asymptotic convergence of multi-fidelity REINFORCE to locally optimal policies in the target environment, and achieves faster finite-sample convergence rates compared to training with high-fidelity data alone. Empirically, we evaluate the MFPG algorithm across a suite of simulated robotics benchmark tasks in scenarios with limited high-fidelity data but abundant off-dynamics, low-fidelity data. With mild to moderate dynamics gaps, MFPG reliably improves the median performance over a standard baseline trained with only high-fidelity data, matching the performance of leading multi-fidelity baselines despite its simplicity and minimal tuning overhead. Under large dynamics gaps, MFPG demonstrates the strongest robustness among the evaluated multi-fidelity approaches. An additional experiment shows that MFPG can remain effective even when low-fidelity environments exhibit reward misspecification. Thus, MFPG not only offers a novel paradigm for efficient sim-to-real transfer, but also provides a principled approach to managing the trade-off between policy performance and data collection costs. Code available at https://github.com/CLeARoboticsLab/MultiFidelityPolicyGradients.
22footnotetext: Equal contribution.
@@ 1 Introduction
Reinforcement learning (RL) algorithms offer significant capabilities in systems that work with unknown, or difficult-to-specify, dynamics and objectives. The flexibility and performance of RL algorithms have led to their adoption in applications as diverse as controlling plasma configurations in nuclear fusion reactors (Degrave et al., 2022), piloting high-speed aerial vehicles (Kaufmann et al., 2023), training reasoning capabilities in large language models (Shao et al., 2024), and searching large design spaces for the automated discovery of new molecules (Bengio et al., 2021; Ghugare et al., 2023). However, in many such applications, datasets must often be gathered from operational systems or from high-fidelity simulations. This requirement acts as a significant barrier to the development and deployment of RL policies: excessive interactions with operational systems are often either infeasible or unsafe, and generating simulated datasets for RL can be prohibitively expensive unless the simulations are both cheap to run and carefully designed to minimize the sim-to-real gap.
@@
More specifically, we present multi-fidelity policy gradients (MFPGs), an RL framework that mixes a small amount of data from the target environment with a control variate formed from a large volume of low-fidelity data to construct an unbiased, variance-reduced estimator for on-policy policy gradients. We further propose a practical algorithm that instantiates the multi-fidelity policy gradient (MFPG) estimator with the classical on-policy algorithm REINFORCE (Williams, 1992).
[…]
$\mathbf{\sum_{t}G_{t}^{h}\log\pi_{\theta}(a_{t}^{h}|s_{t}^{h})}$ $\mathbf{\sum_{t}G_{t}^{l}\log\pi_{\theta}(a_{t}^{l}|s_{t}^{l})}$ policy gradient estimate
@@
Empirically, we evaluate the proposed algorithm on a suite of benchmark robotic control tasks where the high- and low-fidelity environments differ in transition dynamics, and additionally on a task with a misspecified reward function in the low-fidelity environment. In high-fidelity sample-scarce regimes, the results reveal three key insights: i) When the dynamics gap is mild, MFPG can substantially reduce the variance of policy gradient estimates compared to standard variance reduction via state-value baseline subtraction. ii) With mild to moderate dynamics gaps, MFPG reliably outperforms the baseline that only uses high-fidelity data in median performance, matching the strongest multi-fidelity methods despite its algorithmic simplicity and minimal tuning overhead. In more challenging scenarios with large dynamics gaps, MFPG is the most robust among all evaluated multi-fidelity methods, incurring the fewest performance collapses. iii) MFPG can remain effective even in the presence of reward misspecification in the low-fidelity samples.
@@ 2 Related Work
Variance Reduction in RL via control variates. The method of control variates (CVs) is commonly used for variance reduction in Monte Carlo estimation (Owen, 2013). In RL, particularly in policy gradient (PG) methods (Williams, 1992), there is a long history of using CVs to reduce variance and accelerating learning, e.g., subtracting constant reward baselines (Sutton, 1984; Williams, 1988; Dayan, 1991; Williams, 1992) or state value functions (Weaver & Tao, 2001; Greensmith et al., 2004; Peters & Schaal, 2006; Zhao et al., 2011) from Monte Carlo returns. Even though state values are known to not be optimal for variance reduction (Greensmith et al., 2004), they are also central to modern algorithms, e.g., actor-critic methods (Schulman et al., 2017). More recently, additional CVs baselines have been studied, such as vector-form CVs (Zhong et al., 2021), multi-step variance reduction along trajectories (Pankov, 2018; Cheng et al., 2020) and their more general form (Huang & Jiang, 2020), and state-action dependent baselines (Gu et al., 2017; Liu et al., 2018; Grathwohl et al., 2018; Wu et al., 2018), which offer better variance reduction under certain conditions (Tucker et al., 2018). The literature on using CVs for policy gradient s almost exclusively focuses on single-environment settings. In this work, we consider the problem of fusing data from multi-fidelity environments to construct low-variance estimators, cf. Section˜4. Moreover, the proposed multi-fidelity approach can also be readily combined with existing single-fidelity CVs for further improved performance.
@@
Multi-fidelity RL and fine-tuning using high-fidelity data. Multi-fidelity methods for optimization, uncertainty propagation, and inference are well-studied in computational science and engineering, where it is often the case that multiple models for a problem of interest are available (Peherstorfer et al., 2018). Such methods provide principled statistical techniques to manage the computational costs of Monte Carlo simulations, without introducing unwanted biases. However, the adoption of such techniques by RL algorithms is significantly more limited. Two bodies of literature which are closely related to this work are multi-fidelity RL (Cutler et al., 2015) and fine-tuning RL policies using limited target-domain data, e.g., in sim-to-real transfer (Smith et al., 2022). Multi-fidelity RL aims to design training curricula, i.e., decide when to train in each simulation fidelities, in order to train highly-performant policies for the highest-fidelity environment with minimum simulation costs. Typically, training begins in the lowest-fidelity simulators, with proposed mechanisms to decide when to transition to other fidelities based on estimated uncertainty (Suryan et al., 2017; Cutler et al., 2015; Agrawal & McComb, 2024; Ryou et al., 2024; Bhola et al., 2023) in predicted actions, state-action values, dynamics, rewards, and/or information gain (Marco et al., 2017). Similarly, in sim-to-real and transfer learning settings (Taylor & Stone, 2009; Zhao et al., 2020; Tang et al., 2024; Da et al., 2025; Niu et al., 2024), the objective is to train a highly-performant policy for the real world using simulation data, sometimes supplemented with a small amount of real-world data. Many approaches leverage models or value functions trained in simulation (source domain) to bootstrap fine-tuning in the real world (target domain) (Rusu et al., 2017; Arndt et al., 2020; Taylor et al., 2007; Smith et al., 2022), guide real-world exploration (Yin et al., 2025), or bidirectionally align state distributions between simulation and real-world agents to improve sample efficiency. Real-world data can also be used to refine the simulator itself (Abbeel et al., 2006; Chebotar et al., 2019; Ramos et al., 2019). The works above introduce new training paradigms and exploration strategies with data from different domains. In contrast, our approach proposes a new policy gradient estimator that incorporates cross-domain data. The proposed approach can be seamlessly integrated into existing paradigms that involve fine-tuning within target domains.
@@
Off-dynamics RL. A particularly important setting is when multi-fidelity environments exhibit mismatched transition dynamics. Classical approaches such as system identification (Ljung, 1998) and domain randomization (Tobin et al., 2017) address this problem from a dynamics modeling perspective, but typically require either access to a high-fidelity dynamics model or a pre-specified distribution over dynamics parameters. Other works (Desai et al., 2020b; a) instead learn action transformations to compensate for transitions produced by low-fidelity dynamics.
@@
Another line of work tackles the off-dynamics RL problem from a policy adaptation perspective, i.e., modifying policy learning with samples from multiple domains. DARC (Eysenbach et al., ) estimates importance weights for low-fidelity transitions—via a pair of learned classifiers modeling the likelihood ratio under high- and low-fidelity dynamics—and uses these weights to augment the low-fidelity rewards. Guo et al. (2024) extend this idea with generative adversarial imitation learning. Similar in spirit to DARC (Eysenbach et al., ), more recent works estimate the dynamics mismatch by comparing observed low-fidelity transitions with a corresponding high-fidelity prediction—either through explicit target dynamics models to evaluate state-value consistency (Xu et al., 2023) or through latent representation learning (Lyu et al., 2024a). The estimated dynamics gap is then used to filter out unrealistic low-fidelity transitions (Xu et al., 2023) or to augment the low-fidelity rewards by penalizing less likely transitions (Lyu et al., 2024a). A related line of work extends these ideas to offline RL, where high-fidelity data come from a fixed dataset (Niu et al., 2022; Liu et al., 2022; Niu et al., 2025). These methods use learned domain classifiers—originally proposed in DARC (Eysenbach et al., )—to regularize policy learning with low-fidelity samples.
@@ 3 Preliminaries
Our objective is to develop RL algorithms capable of leveraging data generated by multiple distinct environments in order to efficiently learn a policy that achieves high performance in a target environment of interest.
@@
Modeling the multi-fidelity environments. We begin by considering a high-fidelity environment and a low-fidelity environment, each modeled by an MDP. In particular, the high-fidelity environment is an MDP $\mathcal{M}^{h}=(S,A,\Delta_{s_{I}},\gamma,p^{h},R^{h})$, which we assume represents either an accurate simulator of the target environment, or the target environment itself. Here, $S$ is the set of environment states, $A$ is its set of actions, $\Delta_{s_{I}}$ is an initial distribution over states, $\gamma\in[0,1]$ is a discount factor, $p^{h}(s^{\prime}|s,a)$ is the probability of transitioning to state $s^{\prime}$ from state $s$ under action $a$, and $r\sim R^{h}(s,a,s^{\prime})$ defines the probability of observing a particular reward under a given state-action-state triplet. Similarly, the low-fidelity environment is an MDP $\mathcal{M}^{l}=(S,A,$ $\Delta_{s_{I}},\gamma,p^{l},R^{l})$, whose transition dynamics $p^{l}$ and reward function $R^{l}$ differ from those of $\mathcal{M}^{h}$, and do not necessarily accurately represent the target environment. We assume that the cost of generating sample trajectories in $\mathcal{M}^{h}$ is significantly higher than that of generating trajectories in $\mathcal{M}^{l}$, as may be the case if $\mathcal{M}^{h}$ were to represent real-world hardware while $\mathcal{M}^{l}$ were to represent cheap simulations.
@@
Our objective is to learn a stochastic policy $\pi_{\theta}(a|s)$, parameterized by $\theta\in\mathbb{R}^{d}$, that achieves a high expected total reward in the high-fidelity environment $\mathcal{M}^{h}$. Fixing a policy $\pi_{\theta}$ defines a distribution over trajectories in both $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$. We denote trajectories in $\mathcal{M}^{h}$ by $\tau^{h}=s_{0}^{h},a_{0}^{h},r_{0}^{h},s_{1}^{h},a_{1}^{h},r_{1}^{h},\ldots,s_{T}^{h}$, where $s_{0}^{h}\sim\Delta_{s_{I}}$, $a_{t}^{h}\sim\pi_{\theta}(\cdot|s_{t}^{h})$, $s_{t+1}^{h}\sim p^{h}(\cdot|s_{t}^{h},a_{t}^{h})$, and $r_{t}\sim R^{h}(s_{t}^{h},a_{t}^{h},s_{t+1}^{h})$. Similarly, we use $\tau^{l}$ to denote trajectories sampled in $\mathcal{M}^{l}$.
We remark that the proposed method may be extended to integrate data from an arbitrary number of environments, each of which simulates the target environment with a different level of fidelity. One may also define $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$ to have different sets of states and actions, as well as different initial distributions and discount factors. Furthermore, the framework can be extended to infinite-horizon cases. However, neither of these extensions significantly enhances the primary conceptual, theoretical, and experimental contributions of this work, and so we leave them to future research.
@@
Policy gradients. PG algorithms aim to maximize the performance measure $J(\theta)\vcentcolon=\mathbb{E}_{\tau\sim\mathcal{M}(\pi_{\theta})}[R(\tau)]$—the expected total reward along trajectories $\tau$ sampled from environment $\mathcal{M}$ under policy $\pi_{\theta}$. They do so by using stochastic estimates of the policy gradient $\nabla_{\theta}J(\theta)$ to perform gradient ascent directly on the policy parameters (Sutton & Barto, 2018). For example, the REINFORCE algorithm (Williams, 1988; 1992) uses Monte Carlo estimates of $\mathbb{E}_{\tau\sim\mathcal{M}(\pi_{\theta})}[\nabla_{\theta}X_{\tau}^{\pi_{\theta}}]$ to estimate the policy gradient, where $X_{\tau}^{\pi_{\theta}}\vcentcolon=\frac{1}{T}\sum_{t=0}^{T-1}G_{t}\log\pi_{\theta}(a_{t}|s_{t})$ is a random variable defining the contribution of each trajectory $\tau$ to the overall policy gradient. Here, $a_{t}$, $s_{t}$, and $G_{t}=\sum_{k=t}^{T-1}\gamma^{t}r_{t}$, denote the selected action, the state, and the reward-to-go at time $t$ in trajectory $\tau$, respectively.

### v4

@@ Abstract
Many reinforcement learning (RL) algorithms are impractical for deployment in operational systems or for training with computationally expensive high-fidelity simulations, as they require large amounts of data. Meanwhile, low-fidelity simulators—such as reduced-order models, heuristic reward functions, or generative world models—can cheaply provide useful data for RL training, even if they are too coarse for direct sim-to-real transfer. We propose multi-fidelity policy gradients (MFPGs), an RL framework that mixes a small amount of data from the target environment with a control variate formed from a large volume of low-fidelity simulation data to construct an unbiased, variance-reduced estimator for on-policy policy gradients. We instantiate the framework by developing a practical, multi-fidelity variant of the classical REINFORCE algorithm. We show that under standard assumptions, the MFPG estimator guarantees asymptotic convergence of multi-fidelity REINFORCE to locally optimal policies in the target environment, and achieves faster finite-sample convergence rates compared to training with high-fidelity data alone. We evaluate the MFPG algorithm across a suite of simulated robotics benchmark tasks in scenarios with limited high-fidelity data but abundant off-dynamics, low-fidelity data. In our baseline comparisons, for scenarios where low-fidelity data are neutral or beneficial and dynamics gaps are mild to moderate, MFPG is, among the evaluated off-dynamics RL and low-fidelity-only approaches, the only method that consistently achieves statistically significant improvements in mean performance over a baseline trained solely on high-fidelity data. When low-fidelity data become harmful, MFPG exhibits the strongest robustness against performance degradation among the evaluated methods, whereas strong off-dynamics RL methods tend to exploit low-fidelity data aggressively and fail substantially more severely. An additional experiment in which the high- and low-fidelity environments are assigned anti-correlated rewards shows that MFPG can remain effective even when the low-fidelity environment exhibits reward misspecification. Thus, MFPG not only offers a reliable and robust paradigm for exploiting low-fidelity data, e.g., to enable efficient sim-to-real transfer, but also provides a principled approach to managing the trade-off between policy performance and data collection costs.
22footnotetext: Equal contribution.footnotetext: Project website: https://xinjie-liu.github.io/mfpg-rl/
@@ 1 Introduction
Reinforcement learning (RL) algorithms offer significant capabilities in systems that work with unknown, or difficult-to-specify, dynamics and objectives. The flexibility and performance of RL algorithms have led to their adoption in applications as diverse as controlling plasma configurations in nuclear fusion reactors (Degrave et al., 2022), piloting high-speed aerial vehicles (Kaufmann et al., 2023), training reasoning capabilities in large language models (Shao et al., 2024), and searching large design spaces for automated discovery of new molecules (Bengio et al., 2021; Ghugare et al., 2023). However, in many such applications, datasets must be gathered from operational systems or from high-fidelity simulations. This requirement acts as a significant barrier to the development and deployment of RL policies: excessive interactions with operational systems are often either infeasible or unsafe, and generating simulated datasets for RL can be prohibitively expensive unless the simulations are both cheap to run and carefully designed to minimize the sim-to-real gap.
@@
More specifically, we present multi-fidelity policy gradients (MFPGs), an RL framework that mixes a small amount of data from the target environment with a control variate formed from a large volume of low-fidelity data to construct an unbiased, variance-reduced estimator for on-policy policy gradients. We further propose a practical algorithm that instantiates the MFPG estimator with the classical on-policy algorithm REINFORCE (Williams, 1992).
@@
Empirically, we evaluate the proposed algorithm on a suite of benchmark robotic control tasks where the high- and low-fidelity environments differ in transition dynamics, and additionally on a task where the low-fidelity reward is the negative (anti-correlated) version of the high-fidelity reward. We use 20 random seeds per setting and non-parametric bootstrap confidence intervals to assess statistical significance. In high-fidelity sample-scarce regimes, the results reveal three key insights: i) When the dynamics gap is mild, MFPG can substantially reduce the variance of policy gradient estimates compared to standard variance reduction via state-value baseline subtraction. ii) MFPG provides a reliable and robust way to exploit low-fidelity data. In our baseline comparison experiments, for scenarios where low-fidelity data are neutral or beneficial and dynamics gaps are mild to moderate, MFPG is, among the evaluated off-dynamics RL and low-fidelity-only approaches, the only method that consistently and statistically significantly improves the mean performance over a baseline trained solely on high-fidelity data. When low-fidelity data become harmful, MFPG exhibits the strongest robustness against performance degradation among the evaluated off-dynamics RL and low-fidelity-only methods. Strong off-dynamics RL methods tend to exploit low-fidelity data more aggressively, and can achieve high performance at times, but can also fail substantially more severely. iii) MFPG can remain effective even in the presence of reward misspecification in the low-fidelity samples.
@@ 2 Related Work
This section provides a concise overview of three lines of related work: (i) variance reduction in RL via control variates; (ii) multi-fidelity RL and fine-tuning using high-fidelity data for sim-to-real transfer; and (iii) off-dynamics RL.
Variance Reduction in RL via control variates.
The method of control variates (CVs) is commonly used for variance reduction in Monte Carlo estimation (Owen, 2013). In RL, particularly in policy gradient (PG) methods (Williams, 1992), there is a long history of using CVs to reduce variance and accelerating learning, e.g., subtracting constant reward baselines (Sutton, 1984; Williams, 1988; Dayan, 1991; Williams, 1992) or state value functions (Weaver & Tao, 2001; Greensmith et al., 2004; Peters & Schaal, 2006; Zhao et al., 2011) from Monte Carlo returns. Even though state values are known to not be optimal for variance reduction (Greensmith et al., 2004), they are also central to modern algorithms, e.g., actor-critic methods (Schulman et al., 2017). More recently, additional CVs baselines have been studied, such as vector-form CVs (Zhong et al., 2021), multi-step variance reduction along trajectories (Pankov, 2018; Cheng et al., 2020) and their more general form (Huang & Jiang, 2020), and state-action dependent baselines (Gu et al., 2017; Liu et al., 2018; Grathwohl et al., 2018; Wu et al., 2018), which offer better variance reduction under certain conditions (Tucker et al., 2018). The literature on using CVs for policy gradients almost exclusively focuses on single-environment settings. In this work, we consider the problem of fusing data from multi-fidelity environments to construct low-variance estimators, cf. Section˜4. Moreover, the proposed multi-fidelity approach can also be readily combined with existing single-fidelity CVs for further improved performance.
@@
Multi-fidelity RL and fine-tuning using high-fidelity data.
@@
Multi-fidelity methods for optimization, uncertainty propagation, and inference are well-studied in computational science and engineering, where it is often the case that multiple models for a problem of interest are available (Peherstorfer et al., 2018). Such methods provide principled statistical techniques to manage the computational costs of Monte Carlo simulations, without introducing unwanted biases. However, the adoption of such techniques by RL algorithms is significantly more limited. Two bodies of literature which are closely related to this work are multi-fidelity RL (Cutler et al., 2015) and fine-tuning RL policies using limited target-domain data, e.g., in sim-to-real transfer (Smith et al., 2022). Multi-fidelity RL aims to design training curricula, i.e., decide when to train in each simulation fidelities, in order to train highly-performant policies for the highest-fidelity environment with minimum simulation costs. Typically, training begins in the lowest-fidelity simulators, with proposed mechanisms to decide when to transition to other fidelities based on estimated uncertainty (Suryan et al., 2017; Cutler et al., 2015; Agrawal & McComb, 2024; Ryou et al., 2024; Bhola et al., 2023) in predicted actions, state-action values, dynamics, rewards, and/or information gain (Marco et al., 2017). Similarly, in sim-to-real and transfer learning settings (Taylor & Stone, 2009; Zhao et al., 2020; Tang et al., 2024; Da et al., 2025; Niu et al., 2024), the objective is to train a highly-performant policy for the real world using simulation data, sometimes supplemented with a small amount of real-world data. Many approaches leverage models or value functions trained in simulation (source domain) to bootstrap fine-tuning in the real world (target domain) (Rusu et al., 2017; Arndt et al., 2020; Taylor et al., 2007; Smith et al., 2022), guide real-world exploration (Yin et al., 2025), or bidirectionally align state distributions between simulation and real-world agents to improve sample efficiency. Real-world data can also be used to refine the simulator itself (Abbeel et al., 2006; Chebotar et al., 2019; Ramos et al., 2019). The works above introduce new training paradigms and exploration strategies with data from different domains. In contrast, our approach proposes a new policy gradient estimator that incorporates cross-domain data. The proposed approach can be seamlessly integrated into existing paradigms that involve fine-tuning within target domains.
Off-dynamics RL.
@@
A particularly important setting is when multi-fidelity environments exhibit mismatched transition dynamics. Classical approaches such as system identification (Ljung, 1998) and domain randomization (Tobin et al., 2017) address this problem from a dynamics modeling perspective, but typically require either access to a high-fidelity dynamics model or a pre-specified distribution over dynamics parameters. Other works (Desai et al., 2020b; a) instead learn action transformations to compensate for transitions produced by low-fidelity dynamics.
@@
Another line of work tackles the off-dynamics RL problem from a policy adaptation perspective, i.e., modifying policy learning with samples from multiple domains. DARC (Eysenbach et al., ) estimates importance weights for low-fidelity transitions—via a pair of learned classifiers modeling the likelihood ratio under high- and low-fidelity dynamics—and uses these weights to augment the low-fidelity rewards. Guo et al. (2024) extend this idea with generative adversarial imitation learning. Similar in spirit to DARC (Eysenbach et al., ), more recent works estimate the dynamics mismatch by comparing observed low-fidelity transitions with a corresponding high-fidelity prediction—either through explicit target dynamics models to evaluate state-value consistency (Xu et al., 2023) or through latent representation learning (Lyu et al., 2024a). The estimated dynamics gap is then used to filter out unrealistic low-fidelity transitions (Xu et al., 2023) or to augment the low-fidelity rewards by penalizing less likely transitions (Lyu et al., 2024a). Van et al. (2024) introduces a mechanism that re-weights and extrapolates low-fidelity transitions while augmenting rewards, in order to mitigate deficient low-fidelity data support under large dynamics gaps. A related line of work extends these ideas to offline RL, leveraging either online low-fidelity simulation data (Niu et al., 2022; 2025) or a large offline low-fidelity dataset (Liu et al., 2022; Lyu et al., 2025; Guo et al., 2025; Anonymous, 2025) to complement a small high-fidelity dataset.
@@ 3 Preliminaries
Our objective is to develop RL algorithms capable of leveraging data generated by multiple environments in order to efficiently learn a policy that achieves high performance in a target environment of interest.
Modeling multi-fidelity environments.
@@
We consider a high-fidelity environment and a low-fidelity environment, each modeled by a finite-horizon MDP. In particular, the high-fidelity environment is an MDP $\mathcal{M}^{h}=(S,A,\Delta_{s_{I}},\gamma,p^{h},R^{h}{\color[rgb]{0,0,0},T})$, which we assume represents either an accurate simulator of the target environment, or the target environment itself. Here, $S$ is the set of environment states, $A$ is its set of actions, $\Delta_{s_{I}}$ is an initial distribution over states, $\gamma\in[0,1]$ is a discount factor, $p^{h}(s^{\prime}|s,a)$ is the probability of transitioning to state $s^{\prime}$ from state $s$ under action $a$, $r\sim R^{h}(s,a,s^{\prime})$ defines the probability of observing a particular reward under a given state-action-state triplet, and $T$ denotes the finite time horizon length. Similarly, the low-fidelity environment is an MDP $\mathcal{M}^{l}=(S,A,$ $\Delta_{s_{I}},\gamma,p^{l},R^{l}{\color[rgb]{0,0,0},T})$, whose transition dynamics $p^{l}$ and reward function $R^{l}$ differ from those of $\mathcal{M}^{h}$, and do not necessarily accurately represent the target environment.
Assumptions on the multi-fidelity environments.
We assume that the cost of generating sample trajectories in $\mathcal{M}^{h}$ is significantly higher than that of generating trajectories in $\mathcal{M}^{l}$, whose cost we treat as negligible. This may be the case if $\mathcal{M}^{h}$ were to represent real-world hardware while $\mathcal{M}^{l}$ were to represent cheap simulations. In addition, the proposed MFPG approach requires that the low-fidelity simulator can be reset to user-specified states on demand, so that we can construct low-fidelity trajectories whose initial states match those of the high-fidelity rollouts and thereby obtain correlated trajectories.
@@
Our objective is to learn a stochastic policy $\pi_{\theta}(a|s)$, parameterized by $\theta\in\mathbb{R}^{d}$, that achieves a high expected total reward in the high-fidelity environment $\mathcal{M}^{h}$. Fixing a policy $\pi_{\theta}$ defines a distribution over trajectories in both $\mathcal{M}^{h}$ and $\mathcal{M}^{l}$. We denote trajectories in $\mathcal{M}^{h}$ by $\tau^{h}=s_{0}^{h},a_{0}^{h},r_{0}^{h},s_{1}^{h},a_{1}^{h},r_{1}^{h},\ldots,s_{T}^{h}$, where $s_{0}^{h}\sim\Delta_{s_{I}}$, $a_{t}^{h}\sim\pi_{\theta}(\cdot|s_{t}^{h})$, $s_{t+1}^{h}\sim p^{h}(\cdot|s_{t}^{h},a_{t}^{h})$, and $r^{h}_{t}\sim R^{h}(s_{t}^{h},a_{t}^{h},s_{t+1}^{h})$. Similarly, we use $\tau^{l}$ to denote trajectories sampled in $\mathcal{M}^{l}$.
Policy gradients.
@@
PG algorithms aim to maximize the performance measure $J(\theta)\vcentcolon=\mathbb{E}_{\tau\sim\mathcal{M}(\pi_{\theta})}[R(\tau)]$—the expected total reward along trajectories $\tau$ sampled from environment $\mathcal{M}$ under policy $\pi_{\theta}$. They do so by using stochastic estimates of the policy gradient $\nabla_{\theta}J(\theta)$ to perform gradient ascent directly on the policy parameters (Sutton & Barto, 2018). For example, the REINFORCE algorithm (Williams, 1988; 1992) uses Monte Carlo estimates of $\mathbb{E}_{\tau\sim\mathcal{M}(\pi_{\theta})}[\nabla_{\theta}X_{\tau}^{\pi_{\theta}}]$ to estimate the policy gradient, where $X_{\tau}^{\pi_{\theta}}\vcentcolon=\frac{1}{T}\sum_{t=0}^{T-1}G_{t}\log\pi_{\theta}(a_{t}|s_{t})$ is a random variable defining the contribution of each trajectory $\tau$ to the overall policy gradient. Here, $a_{t}$, $s_{t}$, and $G_{t}=\sum_{{\color[rgb]{0,0,0}n}=t}^{T-1}\gamma^{{\color[rgb]{0,0,0}n}-t}r_{{\color[rgb]{0,0,0}n}}$, denote the selected action, the state, and the reward-to-go at time $t$ in trajectory $\tau$, respectively.